OpenStudy (unklerhaukus):

How do i find the particular solution to a second order differential equation \[ay^{\prime\prime}+by^\prime+cy=f(x)\] where \(f(x)\) is a hyperbolic function

OpenStudy (unklerhaukus):

can i use undetermined co-efficients or dosen't this work?

OpenStudy (anonymous):

You can use that, I think..

OpenStudy (unklerhaukus):

\[y^{\prime\prime}-y=\cosh x\] \[y_c^{\prime\prime}-y_c=0\]\[m^2-1=0\]\[m=\pm1\] \[y_c=Ae^x+Be^{-x}\] \[y_p^{\prime\prime}-y+p=\cosh x\]

OpenStudy (anonymous):

Why you have added p here??

OpenStudy (unklerhaukus):

what form can i assume \(y_p\) is of \[y_p=C\cosh x+D\sinh x\]?

OpenStudy (anonymous):

Yes you are right..

OpenStudy (unklerhaukus):

i kinda guessed that , can you link me to a reference ?

OpenStudy (anonymous):

What do you mean?? Put: \[y_p = Ccoshx + Dsinhx\] Take the derivative and find \(y_p''\)..

OpenStudy (anonymous):

\[\huge \color{green}{\frac{d}{dx}sinhx = coshx}\] \[\huge \color{green}{\frac{d}{dx}coshx = sinhx}\]

OpenStudy (zzr0ck3r):

@UnkleRhaukus it is because they all are some form of e^(x) and so we use eulers identity. i dont know if thats what your asking...

OpenStudy (unklerhaukus):

ill try it, i just wasent sure if i had guessed the correct form, i can seam to find it in any table for particular solution

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}-y+p=\cosh x\] \[y_p=C\cosh x+D\sinh x;\qquad y^\prime_p=C\sinh x+D\cosh x;\qquad y^{\prime\prime}_p=C\cosh x+D\sinh x\]

OpenStudy (anonymous):

what is p here??

OpenStudy (unklerhaukus):

ops i dont know, p=0

OpenStudy (unklerhaukus):

no wait i meant \[y_p^{\prime\prime}-y_p=\cosh x\]

OpenStudy (anonymous):

I think we can't put \(Acoshx + Dsinhx\), it will become 0 on putting in that DE..

OpenStudy (unklerhaukus):

\[C\cosh x+D\sinh x-C\cosh x+D\sinh x=\cosh x\] \[0=\cosh x\] so i have made a false assumption,

OpenStudy (anonymous):

We can solve the right hand side for coshx..

OpenStudy (unklerhaukus):

\[\cosh x\neq0\]

OpenStudy (anonymous):

Think about this: May be it will help you: \[\huge \color{blue}{coshx = \frac{e^x + e^{-x}}{2}}\] \[\huge \color{blue}{sinhx = \frac{e^x - e^{-x}}{2}}\]

OpenStudy (anonymous):

If we substitute this in place of coshx then we have particular solution is \(e^x\) and \(e^{-x}\) terms.. Getting or not??

OpenStudy (unklerhaukus):

yeah so there is not a formula for the particular solution if f(x) is hyperbolic, i just have to break the hyperbolic up , i get this but i was hoping there would be a more elegant method, like an assumable form, i guess not

OpenStudy (anonymous):

You could solve it if it is: \[y'' + y = coshx\] By putting: \[y_p = Ccoshx + Dsinhx\]

OpenStudy (unklerhaukus):

@waterineyes \[y_p^{\prime\prime}-y_p=\cosh x\] \[y_p=C\cosh x+D\sinh x;\qquad y^\prime_p=C\sinh x+D\cosh x;\qquad y^{\prime\prime}_p=C\cosh x+D\sinh x\] \[(C\cosh x+D\sinh x)-(C\cosh x+D\sinh x)=\cosh x\] \(0=\cosh x\) False this does not work for some reason

OpenStudy (anonymous):

I said for: \[y'' + y = coshx\] and not: \[y'' - y = coshx\]..

OpenStudy (unklerhaukus):

yeah why is it different with a plus,

OpenStudy (anonymous):

The terms will not get cancel..

OpenStudy (anonymous):

\[2Ccoshx + 2Dsinhx = coshx\]

OpenStudy (unklerhaukus):

yeah i can see that in this case , but in general?

OpenStudy (anonymous):

In general you can use that transformation from coshx to exponential..

OpenStudy (unklerhaukus):

how can i tell which method to use for other question in the set

OpenStudy (anonymous):

Go with exponential..

OpenStudy (unklerhaukus):

ok

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}-y_p=\cosh x\] \[\text{---------}\] \[y_p^{\prime\prime}-y_p=\frac{e^x}2+\frac {e^{-x}}2\] \[y_{p_{1}}=Ce^{x};\qquad y^\prime_{p_1}=Ce^{x};\qquad y^{\prime\prime}_{p_1}=Ce^{x}\] \[Ce^{x}-Ce^{x}-=\frac{e^x}2\]\[0=\frac{e^x}2\] False \[y_{p_{2}}=Ce^{-x};\qquad y^\prime_{p_2}=-Ce^{-x};\qquad y^{\prime\prime}_{p_2}=Ce^{-x}\]\[Ce^{-x}-Ce^{-x}=\frac{e^{-x}}2\]\[0=\frac{e^{-x}}2\] False ???

OpenStudy (foolaroundmath):

I'll again deviate from "guessing" the particular solution, by using the Annihilator Method. \(y'' - y = \cosh{x}\) Let \(y_{p}\) be the particular solution and \(D = dy/dx\) \[(D^{2}-1)y_{p} = \frac{e^{x}+e^{-x}}{2}\] To annihilate \(e^{x}\), we see that operating \((D-1)e^{x} = 0\) and to annihilate \(e^{x}\), we see that operating \((D+1)e^{-x}=0\), THerefore operating (D+1)(D-1) on the equation, we get \[(D^{2}-1)^{2}y_{p}=0 \Rightarrow (D+1)^{2}(D-1)^{2}y_{p} = 0\] so, \(y_{p}\) is of the form \[y_{p}= ae^{x}+bxe^{x}+ce^{-x}+dxe^{-x}\]As the homogenous solution contains \(e^{x}\) and \(e^{-x}\) as solutions, therefore the particular solution is \(\Large y_{p} = bxe^{x}+dxe^{-x}\)

OpenStudy (anonymous):

i got \[y_p=\frac{3xe^x}{8} + \frac{e^x}{16} +\frac {xe^{-x}}{8}-\frac{e^{-x}{8}\] somehow ^ doesn't work on my screen y_p = (3xe^x)/8 + e^x/16 + (xe^-x)/8-e^-x/8 1 bit lemme check wolfram if it's right

OpenStudy (unklerhaukus):

\[y_p=\frac{3xe^x}{8} + \frac{e^x}{16} +\frac {xe^{-x}}{8}-\frac{e^{-x}}{8}\] your just missing a closing brace for the numerator in the last term @nphuongsun93

OpenStudy (anonymous):

o lol =.= *new to latex* but that was wrong.. brb retry

OpenStudy (unklerhaukus):

the answer in the back of the book is \[y=\frac x2\sinh x+Ae^x+Be^{-x}\] so \[y_p=\frac x2\sinh x\]

OpenStudy (foolaroundmath):

@UnkleRhaukus , I do get \(y_{p} = (x \sinh{x})/2\) after using what I have done.

OpenStudy (anonymous):

\[y'' - y = \frac{e^x}{2} + \frac {e^{-x}}{2}\] \[y_p1=\frac{e^x}{2}\] \[y_p = Axe^x\] \[y_p'=A(x+1)e^x\] \[y_p''=A(x+2)e^x\] \[y''-y' = Axe^x+2Ae^x-Axe^x\] \[2Ae^x=\frac{e^x}{2}\] \[A = 1/4\] \[y_p2= Ae^{-x}\]

OpenStudy (anonymous):

zz misclick post too soon D: typing the rest

OpenStudy (unklerhaukus):

\[y_{p_1}^{\prime\prime}-y_{p_1}=\frac{e^x}2\]How do you know to use this form?\[y_{p_1}=Axe^x\] Where does the multiplicative x comes from?

OpenStudy (anonymous):

ignore the last line \[y_p2=Bxe^{-x}\] \[y_p'=-Be^{-x}(x-1)\] \[y_p''=Be^{-x}(x-2)\] \[y''-y=e^{-x}/2\] \[Bxe^{-x}-2Be^{-x}-Bxe^{-x}=e^{-x}/2\] \[B=-1/4\] \[y_p = xe^x/4 - xe^{-x}/4\]

OpenStudy (unklerhaukus):

you have got the solution i want @nphuongsun93 Thank you and well done i just dont understand where the x comes from

OpenStudy (anonymous):

ok your question because the homogenous solution is \[c_1e^x+c_2e^{-x}\] and the coefficients for the particular solution happens to be \[Ae^x and Be^{-x}\] if you leave the same e^x and e^(-x) you will get the wrong answer, so whenever these are the same, just times the coefficients with x.

OpenStudy (anonymous):

same thing when you get the same homogenous solution like \[y''+4y'+4y=0\] \[r=-2,-2\] \[y=c_1e^{-2x}+c_2xe^{-2x}\] sorry i'm not very good with english and wording ;_;

OpenStudy (unklerhaukus):

so it is because \[c_1e^{\color\red x}+c_2e^{\color\red{-x}}\] \[\color\red {x-x}=0\]?

OpenStudy (anonymous):

no you can't add them like that ^ if the homogenous solution is like this \[c_1e^x+c_2e^{-x} \] and the particular solution is like \[Ae^x+Be^{-x}\] the particular solution is gonna get canceled =0 when you uhh.. add them after "derivative" step

OpenStudy (unklerhaukus):

i think that is what i mean

OpenStudy (anonymous):

oh o: ok want me to give an example?

OpenStudy (unklerhaukus):

yes please

OpenStudy (anonymous):

\[y''+y=5cosx\] let's do ^

OpenStudy (unklerhaukus):

\[y_p=A\cos x;\qquad y^\prime_p=-A\sin x ;\qquad y_p^{\prime\prime}=-A\cos x\]

OpenStudy (unklerhaukus):

\[-A\cos x+A\cos x=5\cos x\]\[0=5\]\error

OpenStudy (unklerhaukus):

\[y_p=Ax\cos x;\qquad y^\prime_p=A\cos x-Ax\sin x ;\qquad y_p^{\prime\prime}=-A\sin x-A\sin -Ax\cos x\]

OpenStudy (anonymous):

page got reloaded -_-

OpenStudy (anonymous):

\[y_p = Axcos(x) + Bxsin(x)\]

OpenStudy (unklerhaukus):

whops

OpenStudy (anonymous):

\[y_p''=-2Asinx-Acosx+2Bcosx-2Bxsinx\]

OpenStudy (unklerhaukus):

\[y_p=Ax\cos x+Bx\sin x\]\[ y^\prime_p=A\cos x-Ax\sin x +B\sin x+Bx\cos x\]\[y_p^{\prime\prime}=-A\sin x-A\sin x -Ax\cos x+B\cos x+B\cos x-Bx\sin x\]

OpenStudy (anonymous):

\[y''+y=5cosx\] \[-2Asinx-2Bcosx=5cosx\] \[B=-5/2\]

OpenStudy (anonymous):

\[y_p = \frac{-5}{2}cosx\]

OpenStudy (unklerhaukus):

did i differentiate incorrectly?

OpenStudy (anonymous):

oh no i typed wrong again D: ignore that 2 in the 2B x sin(x). (y'')

OpenStudy (fwizbang):

These are the same kinds of problems as the one we looked at yesterday, where the driving term on the rhs is a solution of the homogeneous diff equation. So, like yesterday, you need to use a different guess.....

OpenStudy (unklerhaukus):

go on,..

OpenStudy (fwizbang):

For the first one(I'm still catching up here) if the equation where \[y \prime \prime-y = \cosh(\lambda x)\] with [\\lambda \neq 1], you'd know the solution was [\ y= A \cosh x + B \sinh x + \cosh x/ (\lambda^2-1) \] yes?

OpenStudy (unklerhaukus):

i did not know that

OpenStudy (fwizbang):

The A and B are the homogeneous solution. For the particular solution, you'd guess that \[y_p= C \cosh(\lambda x)+ D \sinh(\lambda x)\] and then when you plugged this in and solved for C and D, you'd find C=0 and D= 1(lambda^2 -1).

OpenStudy (unklerhaukus):

yeah in my case \(\lambda\) was 1 so \(D=(\lambda-1)=0\)

OpenStudy (anonymous):

Here is the final solution for you to use to check your answer: \[ \begin{array}{c} y(x)=-\frac{1}{4} e^{-x} x+\frac{e^x x}{4}-\frac{e^{-x}}{8}-\frac{e^x}{8}+e^x c_1+e^{-x} c_2 \\ \end{array} \]

OpenStudy (fwizbang):

Right. Here's the "trick" , the solution we're after isn't the particular soution, but really the full solution. In the full solution, A and B are arbitrary constants that we use to set the boundary conditions(initial values) of the problem. In general, they'll depend om lambda. So what wer'e going to do is to choose a slightly different guess for the particular solution by just moving a little of the homogeneous solution into the particular \[y_p = (\cosh \lambda x - \cosh x)/(\lambda^2 -1) \] This is still a solution for lambda not equal to 1, because the extra cosh x term satisfies the homogeneous eq. As lambda goes to 1, this form is indeterminate, so you have to apply L'hopital's rule to evaluate it, and you get \[y_p= xsinh x/2\]

OpenStudy (anonymous):

Or \[ \begin{array}{c} y(x)=-\frac{1}{4} e^{-x} x+\frac{e^x x}{4}+e^x c_1+e^{-x} c_2 \\ \end{array} \]

OpenStudy (anonymous):

or \[ \begin{array}{c} y(x)=\frac{1}{2}x \sinh(x) +e^x c_1+e^{-x} c_2 \\ \end{array} \]

OpenStudy (anonymous):

When choosing the particular solution, you should prevent any part of it to be a solution of the homogeneous solution. This is achieved by multiplying the initial guess by x or x^2

OpenStudy (unklerhaukus):

\[y_p =\lim_{\lambda \rightarrow1} \frac{\cosh \lambda x - \cosh x}{\lambda^2 -1}\rightarrow \frac{0}{0}\] \[\stackrel{\text{l'H}}=\frac{\frac{\text d}{\text d\lambda}\left(\cosh \lambda x - \cosh x\right)}{\frac{\text d}{\text d\lambda}\left(\lambda^2-1\right)}\]\[=\frac{\lambda\sinh \lambda x}{2\lambda}\]\[=\frac{\sinh\lambda x}{2}\]

OpenStudy (unklerhaukus):

im missing an x

OpenStudy (anonymous):

You have to find your particular solution as \( A x e^x + B x e^{-x}\) and find A and B

OpenStudy (unklerhaukus):

\[\stackrel{\text{l'H}}=\frac{\frac{\text d}{\text d\lambda}\left(\cosh \lambda x - \cosh x\right)}{\frac{\text d}{\text d\lambda}\left(\lambda^2-1\right)}=\frac{x\sinh\lambda x}{2\lambda}\]?

OpenStudy (unklerhaukus):

\[y^{\prime\prime}-y=\cosh x\]the complementary solution \[y_c^{\prime\prime}-y_c=0\]\[m^2-1=0\]\[m=\pm1\]\[y_c=Ae^x+Be^{-x}\]

OpenStudy (unklerhaukus):

i see now \[\lim_{\lambda\rightarrow 1}\frac{\frac{\text d}{\text d\lambda}\left(\cosh \lambda x - \cosh x\right)}{\frac{\text d}{\text d\lambda}\left(\lambda^2-1\right)}\]\[=\lim_{\lambda\rightarrow 1}\frac{x\sinh\lambda x}{2\lambda}\]\[=\frac{x\sinh x }2\]

OpenStudy (anonymous):

\[ 2 A e^x-2 B e^{-x}=\frac{1}{2} \left(e^{-x}+e^x\right)\\ A= \frac 1 4 \\ B= -\frac 1 4\\ y_p=x\frac 1 2 \left( \frac {e^x- e^{-x}}{2}\right)=\frac 1 2 x \sinh(x) \]

OpenStudy (anonymous):

Doing it the usual way that I described seems to me a slighter more known approach.

OpenStudy (anonymous):

@satellite73 @lalaly

OpenStudy (unklerhaukus):

@eliassaab i just dont know how i would tell when the particular solutions would be part of the homogeneous solution

OpenStudy (anonymous):

Looking at most of the posts, it seems to me that we are making a very easy problem into a somewhat a bit more difficult.

OpenStudy (unklerhaukus):

this is not a very easy problem

OpenStudy (anonymous):

Use ricatti solving method

OpenStudy (anonymous):

It is very straight forward, believe me. Since \( cosh(x) = \frac 1 2( e^x + e^{-x})\) the initial guess is \( A e^x + B e^{-x} \) these are solutions of the homogenous equation. Multiply by x and you get \( A x e^x + B x e^{-x} \) . Now none of the two in the sum is a solution of the homogeneous equation.

OpenStudy (anonymous):

if you see, for example, e^x or cos(x) or f(x) in the homogenous solution, avoid using those again when you are solving for particular solution by multiplying x to those, for example, xe^x or x cos(x) or x f(x)

OpenStudy (anonymous):

if no one solve this for you, im back in 5/6 hours. PM me and i'll do it for you. I dont have time now Use Ricatti's equation

OpenStudy (fwizbang):

way back at the start, you identified the homogeneous solution as A sinh x +B cosh x, at that point you can look at the fct that appears on the RHS of the DEQ. If one(or a linear combination) of the the fcts in the homogeneous solution appears, then the simple guess won't work and you need to try one of these other methods.

OpenStudy (unklerhaukus):

i think im getting it now! i have to look at my homogenous/complimentary solution, if my particular solution is in the same phase it will resonate this mean a factor of x

OpenStudy (anonymous):

@UnkleRhaukus. What is the verdict: easy or not?

OpenStudy (unklerhaukus):

the whole question for me was where does the factor of x come from, now it makes perfect sense... happy

OpenStudy (unklerhaukus):

justifying the factor of x was not easy.